OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 1/4x + 7 and that pass through the point (-8, 9).

OpenStudy (jim766):

what do you know about perpendicular slopes?

OpenStudy (anonymous):

they're the same

OpenStudy (jim766):

parallel lines slopes are the same.. perpendicular lines slopes are opposite reciprocals (do you know what that means?)

OpenStudy (anonymous):

not really

OpenStudy (jim766):

opposite mean different sign reciprocal means to take a fraction and flip it upside down 3/4 the opp rec would be -4/3 I changed the sign and the flipped it

OpenStudy (jim766):

can you find an equation of a line if you have a slope and a point?

OpenStudy (anonymous):

no

OpenStudy (jim766):

there are 2 ways to do it, one with point slope formuia and the other with y = mx + b they are both about the same amount of work....do you know how your teacher wants you to do it?

OpenStudy (anonymous):

it was just asking to show the parallel and perpendicular way to do it.

OpenStudy (jim766):

right there will be 2 answers one equation will be parallel and the other perpendicular.

OpenStudy (jim766):

lets start with parallel do you recognize y = mx + b?

OpenStudy (anonymous):

yes

OpenStudy (jim766):

what does the m stand for?

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

b is y- intercept

OpenStudy (jim766):

right.... we have a slope what is parallel to the equation in the problem remember (same slope parallel)

OpenStudy (anonymous):

1/4

OpenStudy (jim766):

so we put that into our formula y = mx + b y = 1/4x + b if they gave us a y inercept, we could just plug it in, but they didnt (punks) luckily they did give us an (x,y) do you see it in the problem?

OpenStudy (anonymous):

yes

OpenStudy (jim766):

so lets plug that in 9 = 1/4(8) + b now all you have to do is find the y intercept by solvig for b do you know what you would do first?

OpenStudy (jim766):

opps that was supposed tobe a -8 9 = 1/4(-8) + b

OpenStudy (anonymous):

oh

OpenStudy (jim766):

9 = 1/4(-8) + b 9 = -2 + b +2 +2 11 = b now stick that b in our formula y = 1/4x +11 does this make sense?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yeah